# Advantage Shaping as Surrogate Reward Maximization:

### v1

@@ Appendix A Auxiliary Technical Results
and the conventions $\binom{N}{k}=0$ when $N<k$ and $\binom{N}{0}=1$, we have
@@
hence $\widehat{\rho}_{K}=1-\frac{N^{-}-(K-1)}{N}\,\widehat{f}_{K-1}^{+}$. The boundary conventions cover cases $K=1$ (both weights =1=1) and $M^{-}<K$ (ratio =0=0). ∎
@@
Let $r:=r_{0/2}(y,a)$ and $\rho:=\rho_{\theta}(x,a):=\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}[r_{0/1}(y,a)]$. Define
@@
Appendix B Missing Proofs
B.1 Proof of Equation (14)
@@
B.2 Proof of of Claim 1: Reverse-Engineering $\widetilde{\text{GRPO}}_{K}$
@@
B.3 Proof of Claim 3: Reverse-engineering Skew-R
@@
$G_{\text{skew-R}}=\frac{1}{\sqrt{2\rho(1-\rho)}}\cdot\nabla_{\theta}\rho+\frac{1-2\rho}{\sqrt{2\rho(1-\rho)}}\cdot\nabla_{\theta}\rho\,.$
We know from Corollary 1 that the first term corresponds to a surrogate reward $\arcsin(\sqrt{\rho})$. Analogously, the second term corresponds to $\Omega(\rho)$ for $\Omega$ such that $\Omega^{\prime}(\rho)=\frac{1-2\rho}{\sqrt{2\rho(1-\rho)}}$. Integrating gives (up to a constant) $\Omega(\rho)=\sqrt{\rho(1-\rho)}$ completing the proof of the claim.
B.4 Reward-Level Entropy Regularization
@@ C.1 Reverse-Engineering GRPOK
Integrating the multiplicative factor $(1-\rho)^{K-\frac{3}{2}}\rho^{-\frac{1}{2}}$ over $\rho$ the per-example surrogate reward that gives this gradient is the incomplete beta function with parameters $1/2$ and K1/2K-1/2.
@@
${\mathcal{B}}\Big({\rho};{1/2},{3/2}\Big)=\int_{0}^{\rho}\sqrt{\frac{1-\tau}{\tau}}\mathrm{d}\tau=\int_{0}^{\rho}\frac{1}{\sqrt{2\tau(1-\tau})}\mathrm{d}\tau+\int_{0}^{\rho}\frac{1-2\tau}{\sqrt{2\tau(1-\tau})}\mathrm{d}\tau\,.$
The first integral is $\sqrt{2}\arcsin(\sqrt{\rho})$ and the second $\sqrt{2}\sqrt{\rho(1-\rho)}$, confirming the identity. This means there is no inconsistency in our reverse-engineering strategy. At the same time, the simple three-step forward-engineering recipe operates by substituting population quantities with empirical estimates without accounting for unbiasedness (e.g., via leave-one-out): thus it cannot recover the finite-sample ($N$ not $\gg 1$) version of GRPOK and arrives at skew-R.
@@
C.3 Additional Related Work
In the main body, we compared $\text{GRPO}_{K}$ to its biased variant from [Mah+25] and to $\widetilde{\text{GRPO}}_{K}$ as implemented by [Che+25]. Policy optimization for Pass@K has also been studied by [WK25]. Their approach differs from both ours and [Mah+25] in two important ways.
First, [WK25] do not formulate GRPO-style estimators. The fair comparison is therefore with our $\mathrm{REINFORCE}_{K}$ / $\mathrm{RLOO}_{K}$ estimators from Sec. 3, which directly target the Pass@K objective. Like us, they construct an unbiased estimator of $\nabla_{\theta}\rho_{K,\theta}(x,a)$, but the structure of their estimator is different.
@@ C.3 Additional Related Work
Besides, it is not immediately clear from (35) how to subtract a baseline to reduce variance while maintaining unbiasedness. Instead, going from REINFORCEK to RLOOK in Sec. 3 was rather intuitive. We refer the reader to [WK25, Sec. 4] for more discussion on variance reduction of their method.
For completeness, we can apply our reverse-engineering lens to Eq. (35) and verify that it indeed ascends the Pass@K objective in expectation. Rewriting Eq. (35) more compactly using our notation for conditional gradients $\widehat{\nabla}_{+}$ and $\widehat{\nabla}_{-}={\binom{N^{-}-1}{K-1}}\big/{\binom{N-1}{K-1}}$ and for $\widehat{f}_{K-1}^{-}$, we obtain (for a single $(x,a)$)
@@
Thus, in population, the estimator from [WK25] is indeed maximizing $\rho_{K,\theta}$, consistent with our own derivations. The difference is at finite-samples: (i) how gradient weight is assigned across samples (especially incorrect ones), and (ii) how variance reduction is handled at finite $N$.
Finally, [WK25] explicitly emphasize extending Pass@K-style optimization to non-binary continuous rewards. It would be interesting future work to extend the surrogate reward lens to continuous-reward settings.

### v2

@@ B.1 Auxiliary Technical Results
and the conventions $\binom{N}{k}=0$ when $N<k$ and $\binom{N}{0}=1$, we have for $\widehat{\rho}=\frac{N^{+}}{N}$:
@@
hence $\widehat{\rho}_{K}=1-\frac{N^{-}-(K-1)}{N}\,\widehat{f}_{K-1}^{+}$. The boundary conventions cover cases $K=1$ (both weights =1=1) and $N^{-}<K$ (ratio =0=0). ∎
@@
Let $r:=r_{0/1}(y,a)$ and $\rho:=\rho_{\theta}(x,a):=\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}[r_{0/1}(y,a)]$. With expectations over $y\sim\pi_{\theta}(\cdot|x)$, define
@@
B.2 Proof of Equation (14)
@@
B.3 Proof of of Claim 1: Reverse-Engineering $\widetilde{\text{GRPO}}_{K}$
@@
B.4 Proof of Claim 3: Reverse-engineering Skew-R
@@
$G_{\text{skew-R}}=\frac{1}{2\sqrt{\rho(1-\rho)}}\cdot\nabla_{\theta}\rho+\frac{1-2\rho}{2\sqrt{\rho(1-\rho)}}\cdot\nabla_{\theta}\rho\,.$
We know from Corollary 1 that the first term corresponds to a surrogate reward $\arcsin(\sqrt{\rho})$. Analogously, the second term corresponds to $\Omega(\rho)$ for $\Omega$ such that $\Omega^{\prime}(\rho)=\frac{1-2\rho}{2\sqrt{\rho(1-\rho)}}$. Integrating gives (up to a constant) $\Omega(\rho)=\sqrt{\rho(1-\rho)}$ completing the proof of the claim.
B.5 Reward-Level Entropy Regularization
@@ C.1 Reverse-Engineering GRPOK
Integrating the multiplicative factor $(1-\rho)^{K-\frac{3}{2}}\rho^{-\frac{1}{2}}$ over $\rho$ the per-example surrogate reward that gives this gradient is the incomplete beta function with parameters $1/2$ and K1/2K-1/2:
@@
${\mathcal{B}}\big({\rho};{1/2},{3/2}\big)=\int_{0}^{\rho}\sqrt{\frac{1-\tau}{\tau}}\mathrm{d}\tau=\int_{0}^{\rho}\frac{1}{2\sqrt{\tau(1-\tau})}\mathrm{d}\tau+\int_{0}^{\rho}\frac{1-2\tau}{2\sqrt{\tau(1-\tau})}\mathrm{d}\tau\,.$
The first integral is $\arcsin(\sqrt{\rho})$ and the second $\sqrt{\rho(1-\rho)}$, confirming the identity. This means there is no inconsistency in our reverse-engineering strategy. At the same time, the simple three-step forward-engineering recipe operates by substituting population quantities with empirical estimates without accounting for unbiasedness (e.g., via leave-one-out): thus it cannot recover the finite-sample ($N$ not $\gg 1$) version of GRPOK and arrives at skew-R.
@@
C.3 Comparison to Further Related Work
In the main body, we compared $\text{GRPO}_{K}$ to its biased variant from [sadegh] and to $\widetilde{\text{GRPO}}_{K}$ as implemented by [passk_bydance]. Policy optimization for Pass@K has also been studied by [passk_unbiased_gradient]. Their approach differs from both ours and [sadegh] in two important ways.
First, [passk_unbiased_gradient] do not formulate GRPO-style estimators. The fair comparison is therefore with our $\mathrm{REINFORCE}_{K}$ / $\mathrm{RLOO}_{K}$ estimators from Sec. 3, which directly target the Pass@K objective. Like us, they construct an unbiased estimator of $\nabla_{\theta}\rho_{K,\theta}(x,a)$, but the structure of their estimator is different.
@@ C.3 Comparison to Further Related Work
Besides, it is not immediately clear from (39) how to subtract a baseline to reduce variance while maintaining unbiasedness. Instead, going from REINFORCEK to RLOOK in Sec. 3 was rather intuitive. We refer the reader to [passk_unbiased_gradient, Sec. 4] for more discussion on variance reduction of their method.
For completeness, we can apply our reverse-engineering lens to Eq. (39) and verify that it indeed ascends the Pass@K objective in expectation. Rewriting Eq. (39) more compactly using our notation for conditional gradients $\widehat{\nabla}_{+}$ and $\widehat{\nabla}_{-}$ and for $\widehat{f}_{K-1}^{-}$, we obtain (for a single $(x,a)$)
@@
Thus, in population, the estimator from [passk_unbiased_gradient] is indeed maximizing $\rho_{K,\theta}$, consistent with our own derivations. The difference is at finite-samples: (i) how gradient weight is assigned across samples (especially incorrect ones), and (ii) how variance reduction is handled at finite $N$.
Finally, [passk_unbiased_gradient] explicitly emphasize extending Pass@K-style optimization to non-binary continuous rewards. It would be interesting future work to extend the surrogate reward lens to such reward settings beyond binary.